OpenStudy (anonymous):

integral of 10^x from -1 to 1

OpenStudy (anonymous):

misread sorry i got you...

hartnn (hartnn):

integral of a^x is a^x/log a does this help?

hartnn (hartnn):

\[\int\limits_{}^{}a^xdx=\frac{a^x}{\log a}\]

OpenStudy (anonymous):

Yes that does! How do you know that formula?

OpenStudy (anonymous):

Because \[ \frac{d}{dx} a^x = a^x \log(a) \]

OpenStudy (phi):

the way I remember it is 1) I know how to do \[\int\limits e^{ax} dx\] 2) I know 10 is e to some power a : \(10= e^a \) 3) also \( 10^x = (e^a)^x= e^{ax} \) 4) The "a" to make e^a= 10 is ln(10) (log base e of 10) putting it together \[ \int e^{ln(10)x} dx \] where ln(10) is just a constant